OpenStudy (anonymous):

Write an inequality that models the following situation. The student council is sponsoring a concert as a fund raiser. Tickets are $3 per student and $5 per adult. The student council wants to raise at least $1,000. If 200 students attend, how many adults must attend?

OpenStudy (anonymous):

3x + 5y >= $1,000 The problem says there are 200 students. So replace X with 200. This comes to $600. $600 + 5y >= $1,000 Subtract both sides by $600 and you get 5y >= $400 solve for y and you get 80. So the answer you are looking for is: 600 + 5y >= $1,000 This is your inequality.

OpenStudy (anonymous):

thnx, bro!

OpenStudy (anonymous):

No problem :)